# Erased, But Not Forgotten: Erased Rectified Flow Transformers Still Remain Unsafe Under Concept Attack

### v2

@@ Abstract
Recent advances in text-to-image (T2I) diffusion models have enabled impressive generative capabilities, but they also raise significant safety concerns due to the potential to produce harmful or undesirable content. While concept erasure has been explored as a mitigation strategy, most existing approaches and corresponding attack evaluations are tailored to Stable Diffusion (SD) and exhibit limited effectiveness when transferred to next-generation rectified flow transformers such as Flux. In this work, we present ReFlux, the first concept attack method specifically designed to assess the robustness of concept erasure in the latest rectified flow–based T2I framework. Our approach is motivated by the observation that existing concept erasure techniques, when applied to Flux, fundamentally rely on a phenomenon known as attention localization. Building on this insight, we propose a simple yet effective attack strategy that specifically targets this property. At its core, a reverse-attention optimization strategy is introduced to effectively reactivate suppressed signals while stabilizing attention. This is further reinforced by a velocity-guided dynamic that enhances the robustness of concept reactivation by steering the flow matching process, and a consistency-preserving objective that maintains the global layout and preserves unrelated content. Extensive experiments consistently demonstrate the effectiveness and efficiency of the proposed attack method, establishing a reliable benchmark for evaluating the robustness of concept erasure strategies in rectified flow transformers.
WARNING: This paper contains AI-generated content that may be offensive.
@@ 1 Introduction
In recent years, text-to-image (T2I) generation has achieved remarkable progress driven by the development of diffusion models (Ho et al., 2020; Rombach et al., 2022; Luo et al., 2023) in diverse scenarios. However, these developments raise risks of unintended use. Given inappropriate prompts, diffusion models can often produce undesirable or even harmful outputs such as Not-Safe-For-Work (NSFW) content (Forensics, 2024), as their reliance on large-scale web-scraped training datasets that lack human-level quality control. To mitigate these risks, concept erasure (Gandikota et al., 2023; 2024) has emerged as a practical strategy, selectively suppressing a model’s ability to render specified concepts, styles, or objects.
Figure 1: We present ReFlux, the first concept attack method for next-generation flow-matching T2I framework, requiring only 3.57 MB parameters to restore erased concepts with efficiency and precision, offering a lightweight yet extensible benchmark for erasure robustness. Top row: results of state-of-the-art erasures. Bottom row: ReFlux restores the erased concepts. Blue bars are added for content harmony, yellow framed images are original Flux.1 [dev] generations without erasure.
@@
While concept erasure has been extensively studied and benchmarked in the Stable Diffusion (SD) series (Rombach et al., 2022), which is built upon the DDPM (Ho et al., 2020)/DDIM (Song et al., 2020) + U-Net (Ronneberger et al., 2015) framework, the emergence of Flux (Labs, 2024) introduces a fundamentally different architecture that remains underexplored. Unlike SD, Flux integrates flow matching mechanisms (Liu et al., 2022; Lipman et al., 2022) and transformer-based backbones (Vaswani, 2017), alongside additional Google T5 (Raffel et al., 2020) text encoder and rotary positional encoding (RoPE) (Su et al., 2024) for both textual and pixel embeddings. These architectural distinctions reshape concept representation while also challenging the applicability of existing erasure methods to Flux. Yet, despite the growing adoption of Flux as a next-generation T2I model in research, industry, and safety-critical domains, there lacks a systematic and reliable benchmark to assess the robustness of concept erasure in Flux. This critical gap raises a new question:
@@
Q: Can we employ concept attack to evaluate the robustness of existing concept erasure methods on Flux?
To address this question, we first apply existing attack strategies for concept erasure in SD to Flux, leveraging that both models employ similar denoising-based image generation processes. However, experiments show that these existing attack methods are largely ineffective when applied to Flux. We hypothesize that this discrepancy arises from fundamental architectural differences, which in turn shape the design of erasure and attack techniques for each model. Upon further analysis, we find that Flux encodes a linear relationship between text embeddings and attention maps, fundamentally altering the way information propagates within the model. As a result, all concept erasure methods applied to Flux ultimately converge on a common mechanism—attention localization, where token indices are utilized to precisely identify and suppress targeted content. This suggests that, instead of directly adapting existing attack methods, it is essential to design new attack algorithms that explicitly exploit the attention localization property inherent to Flux.
@@
Building on these findings, we propose ReFlux, a simple yet effective attack method that restores erased concepts by reactivating suppressed attention. Specifically, we firstly introduce an attention reactivation loss to selectively enhance the attention of erased tokens, while maintaining stability through two complementary regularizers: an L2 regularizer that constrains update magnitude to prevent instability, and an entropy regularizer that stabilizes attention distribution to avoid degenerate peaks. To further strengthen reactivation and ensure semantic fidelity, we incorporate an attack-guided velocity loss, which steers flow matching dynamics to amplify erased concepts, and a consistency loss that aligns fine-tuning with authentic generative trajectories, thereby preserving global layout and preventing image distortion. Through lightweight Low-Rank Adaptation (LoRA) (Hu et al., 2021) tuning of text parameters, our method achieves precise reactivation of erased concept while preserving layout fidelity (as shown in Figure 1), revealing that existing defenses suppress only surface-level signals on Flux, while underlying semantics remain erased, but not forgotten.
@@
To the best of our knowledge, this is the first systematic study of concept attack on Flux to evaluate the robustness of erasure methods. Our contributions proceed in three steps:
@@
  • We conduct a systematic analysis of Flux and reveal that existing erasure methods fundamentally depend on an inherent attention localization mechanism. This property also explains why adversarial prompt–based attacks, which are effective in other frameworks, often fail against Flux.
@@
  • We introduce ReFlux, a parameter-efficient fine-tuning method that can accurately restore suppressed concepts in Flux. Our approach combines attention reactivation, principled regularization, velocity-guided optimization, and consistency constraints, forming a complementary set of techniques that enables precise recovery of erased content while preserving the global layout.
@@
  • We conduct extensive experiments across nudity, violence, artistic style, entity, abstraction, relationship, celebrity, and others. Beyond achieving the strongest attack performance, our study reveals persistent conceptual residues beneath erased content, establishing a reliable benchmark for evaluating the robustness of concept erasure on Flux.
@@ 2 Related Work
Rectified flow-based T2I models. T2I diffusion models have made substantial progress recently. Notable contributions include DALL-E series (Ramesh et al., 2021; 2022), GLIDE (Nichol et al., 2021), Imagen (Saharia et al., 2022) and SD series (Rombach et al., 2022; Esser et al., 2024). The latest version is SD 3, which represents a major paradigm shift. It employs a simplified sampling strategy in which the forward noising process is reformulated as rectified flow (Liu et al., 2022), enabling a direct connection between data and noise distributions. Flux (Labs, 2024) is a latest T2I diffusion framework building on recent advances exemplified by SD 3, which adopts a rectified flow transformer backbone and delivers outstanding performance in community ELO evaluations, with notably strong prompt adherence and typography. Recognizing its novelty, future potential, and the existing gap in safety research, we conduct the first systematic study of concept attack in Flux.
Concept erasure. Massive training datasets (e.g., LAION-5B (Schuhmann et al., 2022), COYO-700M (Byeon et al., 2022), and Conceptual 12M (Changpinyo et al., 2021)) may cause T2I models to generate harmful, biased, or sensitive content. To mitigate these issues, concept erasure has emerged as a key research direction in the safety of T2I diffusion models, aiming to selectively suppress a model’s ability to produce specific undesirable concepts. Concept erasure has evolved from traditional attention-editing methods Kumari et al. (2023); Gandikota et al. (2023; 2024); Zhang et al. (2023); Schramowski et al. (2023) to more diverse methods integrating knowledge preservation using strategies like semantic anchor mapping (Lu et al., 2024) and adversarial pruning (Bui et al., 2024; Li et al., 2025; Chavhan et al., 2025). However, the applicability of these methods is often limited by their reliance on the traditional SD 1.5 architecture, which differs essentially from Flux. We note that  Gao et al. (2025) represents the first work proposing a concept erasure method specifically for Flux, inaugurating a novel research direction in the next-generation T2I framework.
Attack evaluations against T2I models. Most existing attack evaluation approaches fundamentally fall under adversarial prompting, which introduces small perturbations to the original prompt. These can be broadly categorized into three categories: The first is traditional projected gradient descent (PGD) optimization (Chin et al., 2024; Wu et al., 2023; Du et al., 2023; Liu et al., 2023; Zhang et al., 2024); The second is training-free approaches (Millière, 2022; Zhuang et al., 2023; Tsai et al., 2024), whose key advantage lies in being an order of magnitude faster than PGD-based methods; The third is Large Language Model (LLM)-based emerging methods (Zhang et al., 2025; Xue et al., 2025), which leverage the reasoning capabilities of LLMs to attack T2I frameworks. While novel and promising, such methods suffer from unstable success rates, API costs, and network latency, limiting their scalability for safety evaluation. Crucially, none of these methods are tailored to the latest rectified flow transformers, and properties of Flux framework may diminish their effectiveness.
3 Challenges in Applying Existing Attack Methods to Flux
Here, we analyze why attack methods effective in SD often fail when applied to Flux. We begin by examining the sentence-level embedding of the T5 encoder (Raffel et al., 2020) and the dual-stream design, which limit the transferability of certain prompt-based strategies. We then study the linear relationship between text embeddings and attention maps, showing how target concepts are localized under erasure. Motivated by this insight, we attempt to reverse the suppression process.
Sentence-level embeddings undermine word-sensitive attacks. With the popularity of adversarial prompt attacks like Ring-A-Bell (Tsai et al., 2024), it seems natural to migrate such techniques to Flux. After all, the core idea is simple yet powerful: derive a concept direction by contrasting prompt pairs. Concretely, given $N$ pairs of prompts $\textbf{P}_{c}^{i},\textbf{P}_{\lnot c}^{i}$ that are semantically similar except for the presence or absence of a target concept $c$ (e.g., “A nude girl” vs. “A clothed girl”), Ring-A-Bell computes their embeddings through the text encoder $f(\cdot)$ and modifies any target prompt P by injecting this concept direction into its embedding:
$\tilde{\textbf{P}}_{\text{cont}}=f(\textbf{P})+\beta\cdot\underbrace{\frac{1}{N}\sum_{i=1}^{N}\Big(f(\textbf{P}_{c}^{i})-f(\textbf{P}_{\lnot c}^{i})\Big)}_{\hat{c}:\,\text{empirical concept direction}}.$ (Ring-A-Bell)
Intuitively, $\hat{c}$ captures the semantic axis corresponding to concept $c$, and the method shifts the embedding of P along this axis to revive suppressed concepts even under strong removal defenses.
However, when migrating to Flux, we encounter a fundamental limitation. Whereas SD employs CLIP, whose embeddings are optimized for word-level similarity, Flux relies on T5, which is designed for sentence-level semantics. As shown in Table 1, we construct the empirical concept vector $\hat{c}$ for “nudity” using the 50 prompt pairs released by Ring-A-Bell, and search for the words in the entire CLIP and T5 tokenizer vocabularies whose embeddings are most similar to this direction. However, the results of T5 are far from rational, confirming that T5’s sentence-level embeddings are context-dependent and fail to capture lexical proximity (see Appendix B for detailed explanation), and thus Ring-A-Bell’s word-sensitive mechanism cannot be directly ported to Flux.
@@ 3 Challenges in Applying Existing Attack Methods to Flux
Computational cost of high-dimensional embeddings. T5 embeddings have a maximum shape of $\mathtt{[256,4096]}$, nearly 18 times larger than CLIP’s $\mathtt{[77,768]}$. This makes PGD-based prompt attacks (Zhang et al., 2024; Chin et al., 2024) prohibitively expensive, as each step requires gradient updates over tens of thousands of vocabulary tokens, with a single optimized prompt often taking over 20 minutes. By contrast, our one-prompt inference requires only the time to generate an image ($\sim$0.5 minutes), making it dramatically more practical on Flux.
@@
Naive attention amplification leads to divergence. An important difference between SD and Flux is that Flux lacks explicit cross attention layers in either dual stream or single stream blocks (see Appendix A for detailed Flux architecture). Yet, through a closer inspection of Flux activations and latent structures, we demonstrate that a linear relationship between text embeddings and attention maps persists, as shown in Figure 2 (a). Specifically, Q-K correlations are established by concatenating textual and pixel embeddings along the last dimension before computing attention weights:
$\textbf{Q}=\mathtt{concat}(\textbf{Q}_{\text{text}},\textbf{Q}_{\text{pixel}}),\quad\textbf{K}=\mathtt{concat}(\textbf{K}_{\text{text}},\textbf{K}_{\text{pixel}}),\quad\textbf{W}_{\mathtt{attn}}=\mathtt{Softmax}(\textbf{QK}^{\top}).$ (1)
@@
Here, $\textbf{W}_{\mathtt{attn}}$ encodes the alignment between prompt tokens and visual features, with the columns indexed by $\mathtt{target\_idx}$ corresponding to specific concept tokens in the input text. As shown in Figure 2 (b), concept erasure methods work by suppressing the attention weights at these target columns, effectively diminishing the model’s focus on the undesired concept (see Appendix C for more illustration). We refer to this phenomenon as attention localization. This naturally inspired us to consider the opposite direction: if erasure succeeds by reducing attention, then attacking the target concept seems straightforward at first: simply amplify its attention. Motivated by this intuition and inspired by prior attention-based editing works  (Zhao et al., 2024; Hertz et al., 2022; Xie et al., 2023), we defined a reverse objective as:
@@
which encourages more attention mass to flow toward the target tokens. However, as illustrated in Figure 2 (c), this straightforward amplification causes the attention maps to divergence, leading to failed generations or severely distorted images. A detailed explanation of this phenomenon is provided in Appendix D. Next, we introduce how we solve the problem by proposing ReFlux.

### v3

@@ Abstract
Recent advances in text-to-image (T2I) diffusion models have brought remarkable generative capabilities, but they also raise serious safety concerns because these models can produce harmful or inappropriate content. Concept erasure has been explored as a way to mitigate such risks, yet most existing approaches and their attack evaluations are built for Stable Diffusion (SD) and show limited effectiveness when applied to next-generation rectified flow transformers (collectively known as the “Flux” family). In this work, we introduce ReFlux, the first concept attack method designed to evaluate the robustness of concept erasure in rectified flow based T2I models. Our approach starts from a key observation: when applied to Flux, current erasure techniques rely heavily on attention localization. Building on this insight, we develop a simple but powerful attack strategy that directly exploits this property. At its core, we use a reverse-attention optimization process to reactivate suppressed signals while stabilizing the attention field. This is reinforced by a velocity-guided dynamic that improves the reliability of concept reactivation by steering the flow matching trajectory, and a consistency-preserving objective that keeps the overall composition intact while protecting unrelated content. Comprehensive experiments demonstrate that our attack achieves both high effectiveness and efficiency, providing a reliable benchmark for assessing the robustness of concept erasure methods in rectified flow transformers. Code is available at https://github.com/nxjiang-jnx/ReFlux.
Figure 1: We present ReFlux, the first concept attack method for next-generation flow matching T2I framework, requiring only 3.57 MB parameters to restore erased concepts with efficiency and precision, offering a lightweight, plug-and-play yet extensible benchmark for erasure robustness. Top row: results of state-of-the-art erased models. Bottom row: ReFlux effectively restores the erased concepts. Blue bars are added for content harmony, yellow framed images are original Flux.1 [dev] generations without erasure.
@@ 1 Introduction
In recent years, text-to-image (T2I) generation has made remarkable strides, powered by the rapid advances of diffusion models [16, 39, 30] across a wide range of applications. Yet, this progress also brings growing ethical concerns and risks of misuse. When prompted with inappropriate text, diffusion models may generate undesirable or even harmful results such as Not-Safe-For-Work (NSFW) imagery [11], largely due to their dependence on massive web-scraped datasets that lack human-level quality control. To address these risks, concept erasure [12, 13] has emerged as a practical approach that aims to selectively disable a model’s ability to depict specific concepts, styles, or objects.
@@
While concept erasure has been widely studied and evaluated on the Stable Diffusion (SD) series [39], which builds upon the DDPM [16]/DDIM [47] framework with a U-Net backbone [40], the emergence of rectified flow–based models [21] (collectively known as the “Flux” family) introduces a fundamentally different design that remains largely unexplored. Unlike SD, Flux combines flow matching mechanisms [26, 24] with transformer-based backbones [50], and incorporates the Google T5 text encoder [36] together with rotary positional encoding (RoPE) [48] for both text and image embeddings. These architectural differences reshape how concepts are represented and also challenge the effectiveness of existing erasure methods when applied to Flux. However, despite Flux’s growing use in research, industry, and safety-critical domains, there is still no systematic or reliable benchmark to evaluate the robustness of concept erasure within this new framework. This gap naturally leads to a key question:
@@
Q: Can concept attacks be used to evaluate the robustness of existing concept erasure methods on Flux?
To explore this question, we begin by applying existing concept attack strategies originally developed for SD to Flux, since both models share a denoising-based image generation process. However, our experiments reveal that these attack methods perform poorly on Flux, showing little ability to reactivate erased concepts. We attribute this gap to fundamental architectural differences that shape how each model handles erasure and attack. Closer analysis shows that Flux still encodes a direct relationship between text embeddings and attention maps, which determines how information flows inside the model. Because of this, all existing erasure techniques on Flux eventually rely on a shared mechanism: attention localization, where token indices are used to pinpoint and suppress target content. This insight suggests that rather than simply reusing SD-based attacks, we need new methods that explicitly take advantage of Flux’s attention localization behavior.
@@
Building on these insights, we propose ReFlux, a simple yet effective attack method that restores erased concepts by reactivating suppressed attention. At the core of our approach is an attention reactivation loss that selectively boosts the attention of erased tokens. To keep the optimization stable, we add two complementary regularizers: an L2 term that limits the size of updates to prevent divergence, and an entropy term that smooths the attention distribution and avoids sharp peaks. To further enhance reactivation while preserving semantic integrity, we introduce an attack-guided velocity loss that steers the flow matching process to strengthen erased concepts, and a consistency loss that aligns the fine-tuning process with the model’s natural generative trajectory, maintaining the overall layout and preventing distortions. With lightweight Low-Rank Adaptation (LoRA) [18] tuning of only 3.57MB text parameters, ReFlux achieves precise concept reactivation while preserving visual structure (as illustrated in Figure 1). This reveals that existing defenses in Flux suppress only surface-level signals, while the deeper semantics remain erased, but not forgotten.
@@
To the best of our knowledge, this work presents the first systematic study of concept attacks on Flux for evaluating the robustness of concept erasure methods. Our main contributions are as follows:
@@
  • Understanding Flux’s internal mechanism. We provide a detailed analysis of Flux and show that existing erasure methods rely heavily on an inherent attention localization behavior inside its MM-DiT blocks. This finding further clarifies why adversarial prompt–based attacks, which are effective in other diffusion frameworks, often fail when transferred to Flux.
@@
  • Developing an effective and lightweight attack. We introduce ReFlux, a parameter-efficient fine-tuning approach that accurately restores suppressed concepts in Flux. It combines attention reactivation, principled regularization, velocity-guided optimization, and consistency constraints into a cohesive design, enabling precise recovery of erased content while maintaining the global layout.
@@
  • Establishing a benchmark for robustness evaluation. We conduct comprehensive experiments across diverse concept categories, including nudity, violence, artistic style, entity, abstraction, relationships, and celebrity. Beyond achieving state-of-the-art attack performance, our results uncover persistent conceptual traces beneath erased content, offering a reliable benchmark for testing the robustness of concept erasure on Flux.
@@ 2 Related Work
Rectified flow-based T2I models. T2I diffusion models have made substantial progress recently. Notable contributions include DALL-E series [38, 37], GLIDE [32], Imagen [41] and SD series [39, 10]. The latest version is SD 3, which represents a major paradigm shift. It employs a simplified sampling strategy in which the forward noising process is reformulated as rectified flow [26], enabling a direct connection between data and noise distributions. Flux [21] is a latest T2I diffusion framework building on recent advances exemplified by SD 3, which adopts a rectified flow transformer backbone and delivers outstanding performance in community ELO evaluations, with notably strong prompt adherence and typography. Recognizing its novelty, future potential, and the existing gap in safety research, we conduct the first systematic study of concept attack in Flux.
Concept erasure. Massive training datasets (e.g., LAION-5B [46], COYO-700M [4], and Conceptual 12M [5]) may cause T2I models to generate harmful, biased, or sensitive content. To mitigate these issues, concept erasure has emerged as a key direction in the safety of T2I diffusion models, aiming to selectively suppress a model’s ability to produce specific undesirable concepts. Concept erasure has evolved from traditional attention-editing methods [20, 12, 13, 55, 44, 22, 1] to more diverse methods integrating knowledge preservation using strategies like semantic anchor mapping [29] and adversarial pruning [3, 23, 6]. However, the applicability of these methods is often limited by their reliance on the traditional SD 1.5 architecture, which differs essentially from Flux. We note that  [14] represents the first work proposing a concept erasure method specifically for Flux, inaugurating a novel research direction in the next-generation T2I framework.
Attack evaluations against T2I models. Most existing attack evaluation approaches fundamentally fall under adversarial prompting, which introduces small perturbations to the original prompt. These methods can be roughly grouped into three categories. The first is traditional projected gradient descent (PGD) optimization [7, 51, 9, 25, 57, 34]; The second is training-free approaches [31, 59, 49], which achieve an order-of-magnitude speedup over PGD methods by avoiding iterative optimization. The third is Large Language Model (LLM)-based emerging methods [54, 53], which leverage the reasoning capabilities of LLMs to attack T2I frameworks. While novel and promising, such methods suffer from unstable success rates, API costs, and network latency, limiting their scalability for safety evaluation. Crucially, none of these methods are tailored to the latest rectified flow transformers, and properties of Flux framework may diminish their effectiveness.
3 Challenges in Applying Existing Attack Methods to Flux
Here, we analyze why attack methods that succeed in SD often fail when transferred to Flux. We start by examining the sentence-level embeddings of the T5 encoder, which limit the transferability of many adversarial prompt–based strategies. Next, we investigate the direct relationship between text embeddings and attention maps, revealing how target concepts become localized during erasure. Guided by this observation, we then design an approach to reverse the suppression process and restore the erased concepts.
Sentence-level embeddings undermine word-sensitive attacks. Given the success of adversarial prompt attacks like Ring-A-Bell [49], it seems natural to extend these techniques to Flux. The underlying idea is simple but powerful: identify a concept direction by contrasting pairs of prompts. Specifically, given $N$ pairs of prompts $\textbf{P}_{c}^{i},\textbf{P}_{\lnot c}^{i}$ that are semantically identical except for the inclusion of a target concept $c$ (e.g., “A nude girl” vs. “A clothed girl”), Ring-A-Bell computes their embeddings using a text encoder $f(\cdot)$ and injects the resulting concept direction into the embedding of a target prompt $\mathbf{P}$:
$\tilde{\textbf{P}}_{\text{cont}}=f(\textbf{P})+\beta\cdot\underbrace{\frac{1}{N}\sum_{i=1}^{N}\Big(f(\textbf{P}_{c}^{i})-f(\textbf{P}_{\lnot c}^{i})\Big)}_{\hat{c}:\,\text{empirical concept direction}}.$ (1)
Intuitively, $\hat{c}$ represents the semantic axis of the target concept $c$, and the method shifts the embedding of $\mathbf{P}$ along this axis to revive suppressed concepts under defenses.
However, when applied to Flux, this strategy encounters a fundamental limitation. SD relies on CLIP embeddings, which are optimized for word-level similarity, whereas Flux uses T5, whose embeddings capture sentence-level semantics. To illustrate this difference, we construct an empirical concept vector $\hat{c}$ for “nudity” using 50 prompt pairs released by Ring-A-Bell and then search for the words in both CLIP and T5 vocabularies whose embeddings are most similar to this direction (see Table 1). While CLIP retrieves semantically related terms, T5 produces irrational results. This confirms that T5’s embeddings are highly context-dependent and fail to model fine-grained lexical proximity (see Appendix B for details). As a result, Ring-A-Bell’s word-sensitive mechanism cannot be directly transferred to Flux.
@@ 3 Challenges in Applying Existing Attack Methods to Flux
Computational cost of high-dimensional embeddings. T5 embeddings have a maximum dimension of $\mathtt{[256,4096]}$, nearly 18 times larger than CLIP’s $\mathtt{[77,768]}$. This makes PGD-based prompt attacks [57, 7] extremely costly, as each step must update gradients across tens of thousands of vocabulary tokens. As a result, optimizing a single prompt can take more than 20 minutes on Flux. In contrast, our one-prompt inference requires only standard image generation time (about 0.5 minutes), as the lightweight fine-tuning for each concept is performed once and reused across all prompts, making it far more practical and efficient for Flux.
@@
Naive attention amplification leads to divergence. A key architectural difference between SD and Flux is that Flux does not include explicit cross-attention layers in either its dual-stream or single-stream blocks (see Appendix A for architectural details). However, a closer examination of Flux’s activations and latent structures reveals that a direct relationship between text embeddings and attention maps still exists, as illustrated in Figure 2(a). Specifically, QK correlations are formed by concatenating textual and pixel embeddings along the last dimension before computing attention weights:
$\textbf{Q}=\mathtt{concat}(\textbf{Q}_{\text{text}},\textbf{Q}_{\text{pixel}}),$ (2)
$\textbf{K}=\mathtt{concat}(\textbf{K}_{\text{text}},\textbf{K}_{\text{pixel}}),$
$\textbf{W}_{\mathtt{attn}}=\mathtt{Softmax}(\textbf{QK}^{\top}).$
@@
Here, $\mathbf{W}_{\mathtt{attn}}$ captures how prompt tokens align with visual features, where columns indexed by $\mathtt{target\_idx}$ correspond to the concept tokens in the input text. As shown in Figure 2(b), concept erasure methods work by suppressing the attention weights at these target columns, effectively pulling the model’s focus away from the unwanted concept (see Appendix C for more examples). We call this effect attention localization. That naturally led us to ask the reverse question: if erasure works by lowering attention, why not try boosting it? Following this intuition, and drawing inspiration from prior attention-based editing works [58, 15, 52], we define a reverse objective as:
@@
which pushes more attention toward the target tokens. However, as shown in Figure 2(c), this simple amplification quickly makes the attention maps blow up, resulting in failed generations or heavily distorted images. A detailed discussion of this behavior can be found in Appendix D. Next, we describe how we address this issue with our proposed method, ReFlux.